1. Tianya reported the test status of the tcpx1 chip. By applying trigger readout during the data readout, progress was made when acquiring data by exposed under the light. Different algorithms were tried, and general functionality was proved, although some unwanted pixels still showed fake hits.
2. Zhang Ying reported the test status of the tcpx2 chip. Different columns and sectors were scanned by scurve method. Test results showed that no obvious gradient can be found in the threshold distribution, implying no obvious IR Drop impact was observed. Some clues showed that the bias value of the IDB current( the bias of the discrimination stage) have direct influence on the analog waveform crosstalk. By decreasing its value, the crosstalk during the calibration seemed to be decreased.
3. Wang Jia reported the test status of the LDO block in the tcpx2 chip. Some attempts showed the LDO can give response when properly set the power input and the load capacitance. However periodic waveform was found, and it is suspected that this was due to the large current load. Further experiment was to be tried.
4. Zhang Liang reported some test status of the DAC blocks in the tcpx2 chip. All the current DACs and voltage DACs showed the same response/problem, that for current DACs, it seems the bit 5th can never go high, and for voltage DACs, the bit7 can never go high. Further check on the design will be made.
5. We are going to do the laser experiment this week.